The fluid levels are measured as hA = 0.2 m, hB = 0.3 m. The oil water interface is at hC = −0.25m, the negative sign indicating that it is below the reservoir base.
Determine the pressure pC at the oil-water interface at point C.
1
Expert's answer
2020-08-19T12:57:37-0400
"P_c=\rho_og(h_a-h_c)+\rho_wg(h_b-h_c)\\P_c=(900)(9.8)(0.2+0.25)+(1000)(9.8)(0.3+0.25)\\P_C=9359\ Pa"